Find the measure of the linear pair
(5x-18)° (4x+45)°

Answer by ikleyn(52780)

.
Find the measure of the linear pair
(5x-18)° (4x+45)°
~~~~~~~~~~~~~~~~~~~~~~~


        A linear pair of angles are always adjacent angles.
        A linear pair of angles always form a straight line.
        They together form a straight angle.
        Two angles forming a linear pair have a common vertex and a common arm.
        Their non-common sides are opposite rays that form a line.


In particular, a linear pair of angles are two supplementary angles, that sum up to 180°.


So, write this equation

    (5x-18) + (4x+45) = 180

and solve it for x

    5x + 4x = 180 + 18 - 45

       9x   =      153

        x   =      153/9 = 17.


So, the angles are  5*17-18 = 67 degrees and  4*17+45 = 113 degrees.    ANSWER

Solved.

--------------------

To learn more about supplementary angles,  read the lessons
    - Angles basics
    - Vertical angles
    - HOW TO solve problems on supplementary, complementary or vertical angles - Examples
    - Solved problems on supplementary and complementary angles
in this site.

You will find there many similar problems, solved with complete explanations.

Consider these lessons as your handbook,  textbook,  guide,  tutorials and  (free of charge)  home teacher.

Learn the subject from there.